# Statistical description and dimension reduction of categorical trajectories with multivariate functional principal components

### v3

@@ 4 Two illustrations in sensory analysis
When participants are instructed to click only on the dominant sensation, i.e. when only one perception can be observed at any given time, the protocol is called Temporal Dominance of Sensation” (TDS, see Pineau et al., (2009) for a reference article). When participants are asked to click on all the sensations they perceive in real time from the same list of descriptors, the protocol is called Temporal Check-All-That-Apply (TCATA, see Castura et al., (2016) for a seminal article). Unlike the TDS protocol, several descriptors (or none at all) can be selected at any given time for TCATA experiments. The resulting data consists of binary trajectories linked to each state, with the value 0 when a descriptor is unclicked at time $t$ and 1 when it is clicked. Thus the difference between TCATA and TDS data is that at each instant $t$, the sum of all binary trajectories can be different from 1 and take values in the set $\{0,1,\ldots,q\}$ for TCATA whereas the sum is always equal to 1 for TDS.
@@
The overall number of experiments is $n=150$ and time has been normalized to be $[0,1]$ for each experiment, resulting, for TDS, in the observed categorical trajectories drawn in Figure 1.
@@ 4.1 Temporal dominance of sensation (TDS) trajectories
Figure 3: Empirical probabilities $\widehat{p}_{j}(t)$, $t\in[0,1]$. The curves are drawn only for the states $j$ whose average probability of occurrence is larger than 5%, that is ti say $\int_{0}^{1}\widehat{p}_{j}(t)dt\geq 0.05$
Most statistical analyses for TDS data are, until now, based on the examination of the evolution over time of the proportion $\widehat{p}_{j}(t)$ of occurrence of each state $j$ (see Pineau et al., (2009)), as shown in Figure 3. The interpretation is then generally based on the succession of the most frequently observed states, and would lead to say, in that experiment, that he first dominant sensation is Lemon, followed by Acid, Sweet and to finish by Salty. As seen below, the signal is much more complex since we are, in this controlled experiment, in presence of three different subpopulations (S04, S06 and S07), which are not detected with the average approach that does not take account of individual temporal variations.
A multivariate FPCA of the trajectories $\mathbf{X}_{i}$, $i=1,\ldots,n$, considering equal weights $w_{j}=1/q$ for $j=1,\ldots,q$ has been performed, giving the same weight to all states $S_{1},\ldots,S_{q}$ (see the Appendix for the the analysis considering weights given in (10)). The first eigenvalue $\lambda_{1}$ represents 23% of the total variance, whereas the second one $\lambda_{2}$ represents 11% of the total variance, the third one 7% and the fourth one 6%. The decrease of the eigenvalues to zero is rather slow (see Figure 4), which is not so surprising since the $\mathbf{X}_{i}$ trajectories are not continuous.
Figure 5: Estimated principal component scores with equal weights $w_{j}$. Different colours are used to distinguish the observations according to the set of gustomer-controlled stimuli.
@@
Table 1: Importance of the different states on each dimension of the MFPCA with equal weights $w_{j}$.
@@
Figure 6: MFPCA with equal weights. First component. Variations around the probability of occurrence related to the first component of the Karhunen-Loève expansion, for the most important categories.
@@
Figure 7: MFPCA with equal weights. Second component. Variations around the probability of occurrence related to the second component of the Karhunen-Loève expansion, for the most important categories.
4.2 Temporal Check All That Apply experiments (TCATA)
The mean trajectories $\widehat{p}_{j}(t)$ for each state $j$ are presented in Figure 8. We can remark that their value is equal to 0 in the time interval $[0,0.2]$. This corresponds to a latency time between the start of the tasting and the first click, corresponding to the occurrence of the first sensation. This latency time is removed in TDS to always have one descriptor selected, but it can be kept here. At the end of the tasting, all the descriptors are automatically unselected, which results in a zero mean value at time $t=1$. The most frequently observed states are Lemon at the beginning of the period, followed by Acid, Sweet, and Salty at the end of the period, as in the TDS evaluations of the signals. As the result of the TCATA experiment can be considered as a set of $q$ binary trajectories, MFPCA can be conducted on them. We have drawn in Figure 12 in the Appendix, the average number of selected states $\sum_{j=1}^{q}\widehat{p}_{j}(t)$ at each instant $t\in[0,1]$. Contrary to TDS experiments in which the number of selected states is always equal to one, we note here a variation over time with a maximum value of 1.5 around time $t=0.6$.
Figure 8: Empirical probability $\widehat{p}_{j}(t)$, $t\in[0,1]$ in TCATA experiments.
The first eigenvalue represents 19% of the total variance, whereas the second one represents 12 % of the total variance (see Figure 13 in the Appendix). As seen in Table 4 in the Appendix, the importance of the different states is a slightly different compared to the analysis of TDS data in Section 4.1. We note that the Basil state appears to be important in both the first and second dimension, allowing to distinguish the S06 experiment, which is the only experiment with basil flavours, from the two other tasting experiments.
The estimated principal component scores are drawn in Figure 14 in the Appendix, whereas the main variations around the mean functions, for the first two dimensions, are drawn in Figure 15 and Figure 16 for the descriptors selected in TDS. The points corresponding to the S04 experiment are characterized by a first principal component score with positive values, related, as seen in Figure 15 to a high probability of occurrence of Lemon in the middle of the tasting, a small probability of occurrence of Acid at the end of the period, and a high probability of occurrence of Sweet and a small probability of occurrence of Salty at the end of the period. Nothing appears on the first component for Acid perception. On the second component, the same points have negative scores, showing a high probability of occurrence of Acid when $t$ is larger than 0.5. The conclusions on the first two axes allow us to recognize the main structure of the original signal (see Figure 2). The same reasoning can be applied to S06 and S07, which are well discriminated by the first two dimensions of the principal components scores.
@@ 5 Concluding remark
The way, presented in this work, of reducing the dimension in a vector space of a panel of categorical trajectories allows for simple interpretation and comparison of individual trajectories. It is also direct to apply that methodology to other experimental protocols such as Temporal Check-All-That-Apply, in which $X_{j}(t)$ and $X_{\ell}(t)$, $\ell\neq j$, can be both equal to one at the same time $t$, whereas this would require to increase considerably the number $q$ of states with continuous time correspondence analysis and Markov processes approaches. It can be easily extended to multivariate categorical trajectories and one could consider simultaneously, in our example, the three experiments made on the same panelists, that is to say $\mathbf{Y}=(Y_{S04},Y_{S06},Y_{S07})$. Finally, this dimension reduction approach can be useful to build predictive models, such as scalar-on-categorical functional data regression models, permitting to use categorical trajectories as explanatory variables in statistical models via the principal components scores. In the example presented in Section 4, it is not difficult to predict, with high accuracy, what is the underlying tasting experiment, among S04, S06 and S07, with a simple linear or quadratic discriminant analysis based on the principal components scores. This decomposition can also be useful to detect outlying trajectories.
Acknowledgement
The Institut de Mathématiques de Bourgogne (UMR UB-CNRS 5584) receives support from the EIPHI Graduate School (contract ANR-17-EURE-0002).
Appendix
@@ Appendix A Analysis of TDS data considering weighted MFPCA
We also consider weights, defined in (10) and equal to $w_{j}=\left(\int_{0}^{1}\widehat{p}_{j}(t)(1-\widehat{p}_{j}(t))dt\right)^{-1}$, that give more importance to the states whose average variance is small, that is to say that are often or very rarely observed (see $w_{p(1-p)}$ in Table 2 for the numerical values). As noted in Remark 2, considering these weights lead to impose the covariance operators $\Gamma_{jj}$ to have the same trace. As seen in Figure 4 on the right, the decrease to zero of the sequence of eigenvalues $\lambda_{r}$ is slower compared to previous analysis with equal weights. We draw in Figure 9 the first two principal components.
@@
The examination of the most important states (see Table 3) in the first and second dimension does exhibit some little difference with the case of equal weights. First Basil appears to be influential in the first and in the second dimension, amplifying the probability of occurrence for negative value of the first component and positive value of the second component (see Figure  10 and Figure 11) even if the weight $w_{j}$ associated to this state is smaller in that unequal weights configuration compared to equal weights MFPCA. This makes it possible to identify, among the S06 tasting experiments, those in which the taste of basil was perceived. Another difference is the presence of Mint in the important variable, particularly on the second and third dimension whereas it was not present at all in the equal weights analysis.
@@
Appendix B Additional tables and figures : TCATA experiments
Figure 12: TCATA experiment. Representation of the mean number $\sum_{j=1}^{D}\widehat{p}_{j}(t)$ of selected states over time.
@@ Appendix B Additional tables and figures : TCATA experiments
Figure 14: Estimated MFCPA scores with TCATA data. Different colours are used to distinguish the observations according to the set of gustometer-controlled stimuli
@@
Figure 16: MFPCA of TCATA data. Second component. Variations around the probability of occurrence related to the second component of the Karhunen-Loève expansion, for the most important categories

### v4

@@ 4 An illustration in sensory analysis
Fifty participants, called panelists, took part in a tasting experiment and were asked to click on the sensation they perceived in real time from a list of descriptors. When participants are instructed to click only on the dominant sensation, i.e. when only one perception can be observed at any given time, the protocol is called “Temporal Dominance of Sensation” (TDS, see Pineau et al.,, 2009 for a reference article). When participants are asked to click on all the sensations they perceive in real time from the same list of descriptors, the protocol is called Temporal Check-All-That-Apply (TCATA, see Castura et al.,, 2016 for a seminal article). Unlike the TDS protocol, several descriptors (or none at all) can be selected at any given time for TCATA experiments. For both cases (TDS and TCATA), the resulting data consists of binary trajectories linked to each state, with the value 0 when a descriptor is unclicked at time $t$ and 1 when it is clicked. The difference between TCATA and TDS data is that at each instant $t$, the sum of all binary trajectories can be different from 1 and take values in the set $\{0,1,\ldots,q\}$ for TCATA whereas the sum is always equal to 1 for TDS. We focus in the following on the TDS analysis (Similar results on TCATA data are reported in the Supplementary Material).
@@
The overall number of experiments is $n=150$ and time has been normalized to be $[0,1]$ for each experiment, resulting, for TDS, in the observed categorical trajectories drawn in Figure 1. The average number of jumps per TDS evaluation is 4, with a minimum of 1 and a maximum of 9.
This dataset is particularly relevant because it relies on controlled data, enabling the assessment of result consistency. In contrast, in sensory analyses using conventional products, the ground truth is unknown, making the evaluation of the statistical power of the method more difficult. Here, the use of known stimuli is crucial for validating the interpretation of the results against theoretical expectations.
@@ 4.1 Temporal dominance of sensation (TDS) trajectories
Figure 3: Empirical probabilities $\widehat{p}_{j}(t)$, $t\in[0,1]$. The curves are drawn only for the states $j$ whose average probability of occurrence is larger than 5%, that is to say $\int_{0}^{1}\widehat{p}_{j}(t)dt\geq 0.05$.
Most statistical analyses for TDS data are, until now, based on the examination of the evolution over time of the proportion $\widehat{p}_{j}(t)$ of occurrence of each state $j$ (see Pineau et al.,, 2009), as shown in Figure 3. The interpretation is then generally based on the succession of the most frequently observed states, and would lead to say, in that experiment, that the first dominant sensation is Lemon, followed by Acid, Sweet and to finish by Salty. As seen below, the signal is much more complex since we are, in this controlled experiment, in presence of three different subpopulations (S04, S06 and S07), which are not detected with the average approach that does not take account of individual temporal variations.
We present in the following Sections analyses based on the MPFCA approach introduced in this paper as well as the results of the continuous time correspondence analysis proposed by Deville, (1982).
4.2 Analysis based on multivariate functional principal components
A multivariate FPCA of the trajectories $\mathbf{X}_{i}$, $i=1,\ldots,n$, considering equal weights $w_{j}=1/q$ for $j=1,\ldots,q$ has been performed, giving the same weight to all states $S_{1},\ldots,S_{q}$ with a discretization of 100 points and a basis of eight B-splines, of order 3, with equispaced interior knots (see the Supplementary Material for the analysis considering the unequal weights scheme given in (10)). The first eigenvalue $\lambda_{1}$ represents 23% of the total variance, whereas the second one $\lambda_{2}$ represents 11% of the total variance, the third one 7% and the fourth one 6%. The decrease of the eigenvalues to zero is rather slow (see Figure 4), which is not so surprising since the $\mathbf{X}_{i}$ trajectories are not continuous.
Figure 4: Proportion of total variance captured by the principal components in $\mathbb{H}$.
Figure 5: Estimated principal component scores. Different gray levels are used to distinguish the observations according to the set of gustomer-controlled stimuli.
@@ 4.2 Analysis based on multivariate functional principal components
Table 1: Importance of the different states on each dimension of the MFPCA.
@@
Figure 6: MFPCA. First component. Variations around the probability of occurrence related to the first component of the Karhunen-Loève expansion, for the most important categories.
@@
Figure 7: MFPCA. Second component. Variations around the probability of occurrence related to the second component of the Karhunen-Loève expansion, for the most important categories.
4.3 Analysis based on continuous time correspondence analysis
An alternative approach in the literature for analyzing TDS data is Categorical Functional Data Analysis (CFDA) described in Remark 1 and which is based on a generalization to continuous time of correspondence analysis. In this study, we employed the R package cfda (Preda et al.,, 2021), considering a basis of eight B-splines (of order 3, with equispaced interior knots) to expand the encoding functions.
One limitation of this approach is that it cannot handle events with a zero probability of occurrence, as illustrated in (9). In practice, this means that it is impossible to compute the encoding function for a specific state at time instants when that state is never selected, an issue observed in Figure 8 and Figure 9 for the state Mint and $t\in[0.5,0.9]$. Moreover, in our opinion, interpretation is more complicated due to the multiplicative decomposition of the individual trajectories, as seen in (9). In contrast, deviations from independence in MFPCA rely on an additive decomposition, as seen is (6), which allows for a more direct interpretation in terms of probability of occurrence.
Figure 8: Continuous time correspondence analysis. Evolution over time of the first encoding functions (see equation (7)). Note that these functions are not defined when the proportion of occurrence is equal to zero, as seen for the state Mint when $t\in[0.5,0.9]$.
Figure 9: Continuous time correspondence analysis. Evolution over time of the first encoding functions (see equation (7)). Note that these functions are not defined when the proportion of occurrence is equal to zero, as seen for the state Mint when $t\in[0.5,0.9]$.
The individual scores computed by the cfda R package are drawn, for the first two dimensions, in Figure 10. It reveals that the first axis clearly distinguishes signal S06 (on the left) from signal S04 (on the right), as it was the case for the first principal component scores of MFPCA (see Figure 5). Indeed, the linear correlation between these two scores is equal to 0.93.
Figure 10: CFDA. Individual scores on the first two components.
The interpretation of the individual scores based on the encoding functions drawn in Figure 8 and  9 is not so easy and, as it was already done in Deville, (1982), we present on the same map in Figure 11, the values of the encoding functions on the first two dimensions, at six different instants in time. In this graphical representation, proximity between points indicates a higher probability of co-occurrence, while greater distance suggests a lower probability, as in correspondence analysis. When examining the large values along the first axis, which primarily correspond to trajectories associated with experiment S04, we observe a proximity between Lemon and Licorice at time $t=0$, followed by Acid at $t=0.2$ and Sweet at $t\in\{0.6,0.8,1\}$. This pattern suggests that signal S04 is initially perceived as Lemon and Licorice, then transitions to Acid, and is ultimately perceived as Sweet toward the end of the evaluation. This interpretation aligns well with the actual signals, with the exception of Licorice, which served as a distractor and was perceived by only a small fraction of the panelists. In contrast, signal S06, associated with negative scores on the first axis, is perceived as Sweet at the beginning of the tasting, which is consistent with the real signals.
As illustrated in Figure 9, the highest values of the encoding functions on the second dimension are predominantly linked to the Mint descriptor, which is positioned in the upper part of the plot. Since Mint was a distractor in this task, its presence confirms that the trajectories associated with this descriptor can be considered as outliers. This observation highlights the sensitivity of CFDA to rare events, much like in correspondence analysis. In the context of CFDA, the second axis appears to primarily capture outlying trajectories, whereas in MFPCA, it distinguishes signal S07 from the other experiments.
Figure 11: Continuous time correspondence analysis. Map of the values of the first two encoding functions for $t\in\{0,0.2,0.4,0.6,0.8,1\}$.
@@ 5 Concluding remark
The way, presented in this work, of reducing the dimension in a vector space of a panel of categorical trajectories allows for simple interpretation and comparison of individual trajectories. That statistical methodology can be useful to describe the trajectories related to other experimental protocols such as Temporal Check-All-That-Apply, in which $X_{j}(t)$ and $X_{\ell}(t)$, $\ell\neq j$, can be both equal to one at the same time $t$, whereas this would require to increase considerably the number $q$ of states with continuous time correspondence analysis and Markov processes approaches. It can be easily extended to multivariate categorical trajectories and one could consider simultaneously, in our example, the three experiments made on the same panelists, that is to say $\mathbf{Y}=(Y_{S04},Y_{S06},Y_{S07})$. Finally, this dimension reduction approach can be useful to build predictive models, such as scalar-on-categorical functional data regression models, permitting to use categorical trajectories as explanatory variables in statistical models via the principal components scores. In the example presented in Section 4, it is not difficult to predict, with high accuracy, what is the underlying tasting experiment, among S04, S06 and S07, with a simple linear or quadratic discriminant analysis based on the principal components scores.
Acknowledgement.
The authors would like to express their sincere gratitude to the two anonymous reviewers for their valuable comments and suggestions, which greatly contributed to improving the initial version of this manuscript. The Institut de Mathématiques de Bourgogne (UMR UBE-CNRS 5584) receives support from the EIPHI Graduate School (contract ANR-17-EURE-0002).
Supplementary Material
@@ Appendix A Analysis of TDS data considering weighted MFPCA
We also consider weights, as those defined in (10) in Remark 2, and equal to
$w_{j}=\left(\int_{0}^{1}\widehat{p}_{j}(t)(1-\widehat{p}_{j}(t))dt\right)^{-1}.$
Introducing such weights in the inner product in $\mathbb{H}$ gives more importance to the states whose average variance is small, that is to say that are often or very rarely observed (see $w_{p(1-p)}$ in Table 2 for the numerical values). As noted in Remark 2, considering these weights lead to impose the covariance operators $\Gamma_{jj}$ to have the same trace. As seen in Figure 12 on the right, the decrease to zero of the sequence of eigenvalues $\lambda_{r}$ is slower compared to previous analysis with equal weights. We draw in Figure 13 the first two principal components.
@@
The examination of the most important states (see Table 3) in the first and second dimension does exhibit some little difference with the case of equal weights. First Basil appears to be influential in the first and in the second dimension, amplifying the probability of occurrence for negative value of the first component and positive value of the second component (see Figure  14 and Figure 15) even if the weight $w_{j}$ associated to this state is smaller in that unequal weights configuration compared to equal weights MFPCA. This makes it possible to identify, among the S06 tasting experiments, those in which the taste of basil was perceived. Another difference is the presence of Mint in the important variable, particularly on the second and third dimension whereas it was not present at all in the equal weights analysis.
@@
Appendix B Analysis of TCATA experiments
The mean trajectories $\widehat{p}_{j}(t)$ for each state $j$ are presented in Figure 16. We can remark that their value is equal to 0 in the time interval $[0,0.2]$. This corresponds to a latency time between the start of the tasting and the first click. This latency time is removed in TDS to always have one descriptor selected, but it can be kept here. At the end of the tasting, all the descriptors are automatically unselected, which results in a zero mean value at time $t=1$. The most frequently observed states are Lemon at the beginning of the period, followed by Acid, Sweet, and Salty at the end of the period, as in the TDS evaluations of the signals. As the result of the TCATA experiment can be considered as a set of $q$ binary trajectories, MFPCA can be conducted on them.
We have drawn in Figure 17 the average number of selected states $\sum_{j=1}^{q}\widehat{p}_{j}(t)$ at each instant $t\in[0,1]$. Contrary to TDS experiments in which the number of selected states is always equal to one, we note here a variation over time, with a maximum value of 1.5 selected states around time 0.6. In that sample, the average number of jumps per TCATA trajectory is 7, with a minimum of 1 and a maximum of 21.
The first eigenvalue represents 19% of the total variance, whereas the second one represents 12 % of the total variance (see Figure 18). As seen in Table 4 the importance of the different states is a slightly different compared to the analysis of TDS data. We note that the Basil state appears to be important in both the first and second dimension, allowing to distinguish the S06 experiment, the only experiment with basil flavours, from the two other tasting experiments.
Figure 16: TCATA experiment. Empirical probabilities of occurrence $\widehat{p}_{j}(t)$, for the most frequently selected states.
The estimated principal component scores are drawn in Figure 19, whereas the main variations around the mean functions, for the first two dimensions, are drawn in Figure 20 and Figure 21 for the same states as those selected in the TDS sample. The points corresponding to the S04 experiment are characterized by a first principal component taking positive values, related, as seen in Figure 20 to a high probability of occurrence of Lemon in the middle of the tasting, a small probability of occurrence of Acid at the end of the period, and a high probability of occurrence of Sweet and a small probability of occurrence of Salty at the end of the period. Nothing appears on the first component for Acid perception. On the second component, the same points have negative scores, showing a high probability of Acid occurring after 0.5. The conclusions on the first two axes allow us to recognize the main structure of the original signal (see Figure 2). The same reasoning can be applied to S06 and S07, which are well discriminated by the first two dimensions of the 1-2 map.
Figure 17: TCATA experiment. Mean number $\sum_{j=1}^{D}\widehat{p}_{j}(t)$ of selected states over time.
@@ Appendix B Analysis of TCATA experiments
Figure 19: Estimated MFCPA scores with TCATA data. Different levels of gray are used to distinguish the observations according to the set of gustometer-controlled stimuli.
@@
Figure 21: MFPCA of TCATA data. Second component. Variations around the probability of occurrence related to the second component of the Karhunen-Loève expansion, for the most important categories.
Appendix C A simple example of process taking values in $D([0,T],\mathbb{R})$
We present in this Section a simple example of stochastic process whose trajectories are not continuous and belong to $D([0,T],\mathbb{R})$. The stochastic process is continuous in probability, with mean trajectory that is a smooth function and a covariance function that is not differentiable in the neighborhood of the diagonal. This example is derived from the well known telegraph process. Define
$X(t)=\dfrac{(-1)^{N(t)}+1}{2},$
where $N(t)$ is a Poisson process defined on $[0,T]$, with intensity $\lambda>0$. For every $t\in[0,T]$, $X(t)\in\{0,1\}$. It is well known that the trajectories of $N(t)$ are not continuous, and the number of jumps (discontinuity points) on $[0,T]$ follows a Poisson distribution with parameter $\lambda T$. Thus, for all $\omega\in\Omega$, the number of jumps of $X(.,\omega)$ is finite and $X(.,\omega)\in D([0,T],\mathbb{R})$.
Furthermore, the trajectories of $X(t)$ are continuous in probability. Indeed, we have
$X(t+h)-X(t)=\frac{(-1)^{N(t+h)}-(-1)^{N(t)}}{2},$
and $\mathbb{P}[X(t)\neq X(t+h)]$ is equal to the probability that N(t+h)N(t)N(t+h)-N(t) is an odd number, which is equal to
$\mathbb{P}[X(t)\neq X(t+h)]$ $=\sum_{k=0}^{\infty}\mathbb{P}(N(t+h)-N(t)=2k+1)$
$=\sum_{k=0}^{\infty}\dfrac{e^{-\lambda h}(\lambda h)^{2k+1}}{(2k+1)!}$
$=e^{-\lambda h}\sinh(\lambda h).$
For sufficiently small $h$, we get with a Taylor expansion that $e^{-\lambda h}\sinh(\lambda h)=\lambda h+o(h)$ and $\lim_{h\to 0}\mathbb{P}[X(t)\neq X(t+h)]=0$, showing that for all $t\in[0,T]$, the process $X(t)$ is continuous in probability. Furthermore it is possible to get an explicit expression for its expectation, whose derivatives of any order are well defined. First note that
$\mathbb{E}[(-1)^{N(t)}]$ $=\sum_{k=0}^{\infty}(-1)^{k}\mathbb{P}[N(t)=k]$
$=\exp(-2\lambda t),$
so that
$\mathbb{E}[X(t)]=\frac{\mathbb{E}[(-1)^{N(t)}]+1}{2}=\frac{e^{-2\lambda t}+1}{2}.$
Using the fact that, for $t\geq s$, N(t)N(s)N(t)-N(s) has a Poisson distribution, with parameter $\lambda(t-s)$, we get that
$\text{Cov}(X(s),X(t))=\frac{e^{-2\lambda|t-s|}-e^{-2\lambda(s+t)}}{4}.$
The partial derivatives of $\text{Cov}(X(s),X(t))$ are not defined on the diagonal $(t,t)$ and we can deduce (see Loève,, 1978, Chapter 11) that the stochastic process $X(t)$ is not mean square differentiable. Indeed, we can check that the process $X(t)$ is not mean square differentiable. We have
$\mathbb{E}\left[\left(\frac{X(t+h)-X(t)}{h}\right)^{2}\right]$ $=\frac{1}{4h^{2}}\mathbb{E}\left[\left((-1)^{N(t+h)}-(-1)^{N(t)}\right)^{2}\right]$
$=\frac{e^{-\lambda h}\sinh(\lambda h)}{4h^{2}}.$
Taking the limit as $h\to 0$, we get
$\lim_{h\to 0}\mathbb{E}\left[\left(\frac{X(t+h)-X(t)}{h}\right)^{2}\right]=\lim_{h\to 0}\frac{\lambda}{4h}=+\infty,$
and conclude that the stochastic process $X(t)$ is not mean square differentiable.
Appendix D Noisy measurements of jump times
The jump times of a trajectory $\mathbf{X}_{i}\in D([0,1],\mathbb{R}^{q})$ are defined as follows
$J_{i}=\{t\in[0,1];\mathbf{X}(t-)\neq\mathbf{X}(t)\}$ (17)
The cardinal of $J_{i}$ is finite, say $n_{i}$ and we denote by $\{t_{i1}<\ldots<t_{in_{i}}\}$ its elements ordered in ascending order. As suggested by an anonymous referee, one could consider that the observed jump times are corrupted by noise, so that we effectively observe the jumps at the noisy time instants
$\widetilde{t}_{ik}=t_{ik}+\epsilon_{ik},\quad k=1,\ldots,n_{i},$
with the following correction on the edges of the interval, $\widetilde{t}_{ik}=0$ if $t_{ik}+\epsilon_{ik}<0$ and $\widetilde{t}_{ik}=1$ if $t_{ik}+\epsilon_{ik}>1$.
We have evaluated the influence of such noisy observations on the results of our analysis, considering iid noise $\epsilon_{ik}$, with uniform distribution on [a,a][-a,a] with a = 0.05, 0.1 and 0.2. Noise was added on the jump instants of TDS data presented in Figure 1 and MFPCA, using the same parameters as previously (8 B-splines of order 3, with equispaced knots and time discretized in 100 equispaced points in $[0,1]$). A hundred simulations were performed for each scenario.
The average absolute correlations between the individual principal component scores of the first two components and those of the original (noise-free) first components were computed and are presented in Figure 22. As expected, when the noise variance is small ($a=0.05$), the correlations remain very close to one between the original scores and those based on noisy data. As it increases ($a=0.1$ and $a=0.2$), the correlation decreases, unsurprisingly, with a more pronounced reduction for the second component than for the first.
Figure 22: Distribution of the correlation coefficient between original and noisy first two principal components with time jumps corrupted by an additive noise with uniform distribution over [a,a][-a,a], for $a\in\{0.05,0.1,0.2\}$.